Through (5,-3) slope -1
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

equation of line, if that is desired.
point slope y-y1=m(x-x1), m slope and (x1, y1) point)
y+3=-1(x-5)
y=-x+5-3
y=-x+2

Answer by Alan3354(69443)

What about it?
Do you have a question?